what is angle of dispersion in a prism

Grade:10

4 Answers

Akash Kumar Dutta
11 years ago

i think it is same as the apex angle of the prism or the angle of the prism.

since it decides the angle of deviation as i+e=D+A(angle of the prism),D=deviation.

Rishabh Maggirwar
11 years ago

Did not quite understand the question.

Angle of dispersion is different for different wavelenghts of light.

the most is for violet and least for red.

Akash Kumar Dutta
11 years ago

@ Rishabh
sorry to comment but thats not the angle of dispersion...thats the angle of deviation.
hope i made it correct.

tejas naresh patel
11 years ago

same as the apex angle of the prism or the angle of the prism.